OpenStudy (anonymous):

Does anyone know what to do? 1/ex=2

OpenStudy (anonymous):

\[\frac{ 1 }{ ex }=2 \]

OpenStudy (aum):

\(\Large \frac{1}{e^x} = 2\) ?

OpenStudy (anonymous):

yes

OpenStudy (aum):

\(\Large \frac{1}{e^x} = 2\) \(\Large e^{-x} = 2\) Take natural logarithm on both sides: \(\large \ln(e^{-x}) = \ln(2)\) \(\large -x * \ln(e) = \ln(2)\) \(\large -x * 1 = \ln(2)\) \(\large x = -\ln(2)\)

OpenStudy (aum):

OR \(\large x = -\ln(2) = \ln(2)^{-1} = \ln(1/2) = \ln(0.5)\)

OpenStudy (anonymous):

ok ao since it's a fraction we need to get rid of the fraction by converting it to e(exponent -x)

OpenStudy (anonymous):

Thank you very much for your help! :)

OpenStudy (aum):

It becomes easier to solve. You can also take natural logarithm of the first line: ln(1/e^x) = ln(2) ln(A/B) = ln(A) - ln(B) ln(1/e^x) = ln(1) - ln(e^x) = ln(2) ln(1) = 0 ln(e^x) = x. So 0 - x = ln(2) x = -ln(2)

OpenStudy (anonymous):

the answer i got is x=-0.6931471806

OpenStudy (aum):

correct. Round it to whatever decimal places they tell you to. But if none specified, I will leave the answer as x = ln(0.5) or x = -ln(2).

OpenStudy (anonymous):

oh ok the first one was easier to understand than the second one. Thank you very much!

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

I have another question, just so I can see if I understood the problem. I have another problem \[\ln (2x)=7\]\[x \times \ln (2)=\ln (7)\]\[x \times0.69=1.94\]\[\frac{ 0.69 }{ 0.69 }x=\frac{ 1.94 }{ 0.69 }\]\[x=2.811\] am i right or wrong?

OpenStudy (aum):

\(\ln(AB) = \ln(A) + \ln(B)\)

OpenStudy (aum):

\(\ln(2x) = 7\) implies \(2x = e^7; ~~x = \frac 12 e^7\)

OpenStudy (aum):

\(\ln(2x) \ne x \times \ln(2)\) \(\ln(2x) = \ln(2) + \ln(x)\)

OpenStudy (aum):

Or is it \(2^x\) and not \(2x\) ?

OpenStudy (anonymous):

not exponent, it's 2x

OpenStudy (aum):

Ok, then what I said above is correct. \(\ln(2x) = \ln(2) + \ln(x)\) \(\ln(2^x) = x \times \ln(2) \)

OpenStudy (anonymous):

thank you !

OpenStudy (aum):

you are welcome.